LeAnn wants to giftwrap a present she got for her little brother. How many square inches of giftwrap will be needed to cover a box that is 5in x 7in x 3in?


Answer by amalm06(224)

Let length=l=5,width=w=7,height=h=3
The box is 3-dimensional, with 6 faces. So we have three possible combinations, and for each combination we should multiply by 2 to account for a parallel face
(5)(7)(2)=70
(5)(3)(2)=30
(7)(3)(2)=42
70+30+42=142 sq in (Answer)